# Stable dark matter from Pauli blocking in the degenerate fermion background with Quantum Field Theory

### v1

@@ 2.2 With Cosmic Expansion
In the expanding Universe, the magnitude of a physical momentum $k(t)\equiv|\vec{\bf k}(t)|$ redshifts inversely proportional to the scale factor in the metric of the spacetime. That is also true for the number density of the scalar field, in this case $n_{\phi}\propto a^{-3}$. Under these change of background, the scalar field $\phi$ can decay and fill the Fermi surface. However, the redshift of the momentum make the Fermi surface empty and allows the decay of $\phi$. This happens continuously until the number density of $\phi$ becomes smaller than the number density in the Fermi surface of momentum $p_{\psi}$, when Pauli blocking is not effective any more.
@@
$\begin{split}k(t_{p},t)=\frac{a(t_{p})}{a(t)}p_{\psi},\end{split}$ (2.23)
where $t_{p}$ is the production time of fermion $\psi$ with momentum $p_{\psi}$. In matter-dominated Universe with $a(t)^{3}\propto t^{2}$, $k(t)$ has the relation to the time as
$\begin{split}t_{p}=t\left(\frac{|\vec{\bf k}(t)|}{p_{\psi}}\right)^{3/2}.\end{split}$ (2.24)
For a given comoving momentum, we can write the Boltzmann equation Eq. (2.10) in terms of a total derivative about time as [11]
$\begin{split}\frac{df_{\psi}(t_{p},t)}{dt}=f_{\psi}^{\rm coll}(t_{p},t)\simeq\Gamma_{\phi}n_{\phi}(t)\left(\frac{\pi^{2}}{p_{\psi}^{2}}\right)\frac{\delta(t-t_{p})}{p_{\psi}H(t_{p})}\left[1-2f_{\psi}(\vec{\bf k}(t),t)\right],\end{split}$ (2.25)
@@
$\begin{split}\delta(k(t)-p_{\psi})=\frac{\delta(t-t_{p})}{p_{\psi}H(t_{p})}.\end{split}$ (2.26)
@@
$\begin{split}f_{\psi}(t_{p},t)=\frac{1}{2}\left(1-e^{-\bar{f}(t_{p})}\right)\theta(p_{\psi}-k(t)),\end{split}$ (2.27)
@@
$\begin{split}\bar{f}(t_{p})=\frac{\pi^{2}\Gamma_{\phi}}{p_{\psi}^{3}}\frac{n_{\phi}(t_{p})}{H(t_{p})}.\end{split}$ (2.28)
To fill the Fermi surface of the momentum $k=p_{\psi}$, i.e. $f_{\psi}=1/2$, it is needed that $\bar{f}(t_{p})\gg 1$, or the cosmic time must be large enough to satisfy
$\begin{split}t_{p}\simeq\frac{2}{3H(t_{p})}\gg\left(\frac{n_{\phi}(t_{p})}{p_{\psi}^{3}/\pi^{2}}\right)^{-1}\Gamma_{\phi}^{-1},\end{split}$ (2.29)
which is consistent with that in the non-expanding case in Eq. (2.22).
In Fig. 2 (left), we show the time evolution of the number density of scalar dark matter (solid line) and fermion field (dashed line) for different initial number density of scalar field with $10^{-30}\,\mathrm{GeV}$ (orange) and $n_{\phi}(t_{i})=10^{-25}\,\mathrm{GeV}$ (blue), respectively. For comparison, we also show the case of scalar decay without Pauli blocking (dotted line) for each case.
When $n_{\phi}(t)\gg p_{\psi}^{3}/2\pi^{2}$, the fermions produced from the decay fills the fermi surface $k=p_{\psi}$ quickly with $f_{k}=1/2$ and the number density of them is
$\begin{split}n_{\psi}(t)=&g\int^{p_{\psi}}\frac{d^{3}k}{(2\pi)^{3}}f_{\psi}(k)=\frac{p_{\psi}^{3}}{6\pi^{2}},\end{split}$ (2.30)
@@
$\begin{split}\rho_{\psi}=\frac{p_{\psi}^{4}}{8\pi^{2}}.\end{split}$ (2.31)
@@
$\begin{split}\frac{d}{dt}(\rho_{\phi}+\rho_{\psi})+3H\rho_{\phi}+4H\rho_{\psi}=0.\end{split}$ (2.32)
@@
$\begin{split}\rho_{\phi}(t)=-\frac{p_{\psi}^{4}}{6\pi^{2}}+\left(\rho_{\phi}(t_{i})+\frac{p_{\psi}^{4}}{6\pi^{2}}\right)\left(\frac{a(t)}{a(t_{i})}\right)^{-3},\end{split}$ (2.33)
with $t\gg t_{i}$.
Once the number density of $\phi$ is small, i.e.,
$\begin{split}n_{\phi}(t)\lesssim\frac{p_{\psi}^{3}}{6\pi^{2}},\end{split}$ (2.34)
$\phi$ rapidly decays and $n_{\psi}$ begins to decrease only due to expansion of the Universe.
@@
$\begin{split}f_{\psi}(k,t_{0})=f_{\psi}\left(p_{\psi},t_{0}\left(\frac{k}{p_{\psi}}\right)^{3/2}\right).\end{split}$ (2.35)
@@
Figure 2: The time evolution of the number density of scalar (solid) and fermion (dashed) (left) and the distribution function of the produced fermions at momentum $p_{\psi}$ (right). Here, we used $m_{\phi}=10^{-3}\,{\rm eV},\ f_{\psi}(p_{\psi},t_{i})=0$ and $\Gamma_{\phi}=10^{-37}\,{\rm GeV}$, with $t_{0}=1/\Gamma_{\phi}$. The orange and blue lines correspond to the different initial number density, $n_{\phi}(t_{i})=10^{-30},10^{-25}\,\mathrm{GeV}$, respectively. The lifetime becomes longer for larger number density due to the Pauli blocking.
@@ 3 Particle Production with the Quantum Field Theory
$\begin{split}\psi(x)&=\sum_{s}\int\frac{d^{3}p}{(2\pi)^{3}}\frac{1}{\sqrt{2\omega_{p}}}\left(a_{\vec{p}}^{s}(t)u_{p}^{s}e^{-i\vec{p}\cdot\vec{x}}+b_{\vec{p}}^{s\dagger}(t)v_{\vec{p}}^{s}e^{i\vec{p}\cdot\vec{x}}\right),\\ \bar{\psi}(x)&=\sum_{s}\int\frac{d^{3}p}{(2\pi)^{3}}\frac{1}{\sqrt{2\omega_{p}}}\left(a_{\vec{p}}^{s\dagger}(t)\bar{u}_{p}^{s}e^{i\vec{p}\cdot\vec{x}}+b_{\vec{p}}^{s}(t)\bar{v}_{\vec{p}}^{s}e^{-i\vec{p}\cdot\vec{x}}\right),\end{split}$ (3.1)
where the creation and annihilation operators satisfy the anti-communication relations as
$\begin{split}\{a^{r}_{\vec{\bf k}},a^{s\dagger}_{\vec{\bf p}}\}=\{b^{r}_{\vec{\bf k}},b^{s\dagger}_{\vec{\bf p}}\}=(2\pi)^{3}\delta^{(3)}({\vec{\bf k}}-{\vec{\bf p}})\delta^{rs},\end{split}$ (3.2)
@@
$\begin{split}H_{\rm free}=\int\frac{d^{3}p}{(2\pi)^{3}}\sum_{s}\omega_{\vec{\bf p}}(a^{s\dagger}_{\vec{\bf p}}a^{s}_{\vec{\bf p}}+b^{s\dagger}_{\vec{\bf p}}b^{s}_{\vec{\bf p}}).\end{split}$ (3.3)
The interaction Hamiltonian for Yukawa interaction Eq. (2.1) is
$\begin{split}H_{\rm int}&(t)=\lambda\int d^{3}x\phi(x)\overline{\psi}(x)\psi(x)\\ =&\lambda\phi(t)\int\frac{d^{3}p}{(2\pi)^{3}}\sum_{s,s^{\prime}}\frac{1}{2\omega_{p}}\left[a^{s\dagger}_{p}a^{s^{\prime}}_{p}\bar{u}^{s}_{p}u^{s^{\prime}}_{p}+b^{s\dagger}_{p}b^{s^{\prime}}_{p}\bar{v}^{s}_{p}u^{v^{\prime}}_{p}+b^{s^{\prime}}_{p}a^{s}_{-p}\bar{v}^{s^{\prime}}_{p}u^{s}_{-p}+a^{s^{\prime}\dagger}_{p}b^{s\dagger}_{-p}\bar{u}^{s^{\prime}}_{p}v^{s}_{-p}\right].\end{split}$ (3.4)
In the second line, we assumed that the decaying scalar field is homogeneous and classical.
We introduce the spin-averaged distribution function of $\psi$ and $\bar{\psi}$ as,
$\begin{split}f_{\psi,\vec{k}}(t)\equiv\frac{1}{2}\sum_{r}\frac{1}{V}\left\langle a_{\vec{k}}^{\dagger r}a_{\vec{k}}^{r}\right\rangle,\quad f_{\bar{\psi},\vec{k}}(t)\equiv\frac{1}{2}\sum_{r}\frac{1}{V}\left\langle b_{\vec{k}}^{\dagger r}b_{\vec{k}}^{r}\right\rangle,\end{split}$ (3.5)
where $V$ is spatial volume.
The expectation value of the operator $\mathcal{O}$ is defined using a density matrix $\rho$ as
$\begin{split}\expectationvalue{\mathcal{O}}\equiv{\rm Tr}(\rho\mathcal{O}),\end{split}$ (3.6)
and its time derivative is given by
$\begin{split}\frac{d}{dt}\expectationvalue{\mathcal{O}}=&-i\expectationvalue{\commutator{\mathcal{O}}{H_{\rm int}}}+\dot{\expectationvalue{\mathcal{O}}},\end{split}$ (3.7)
where
$\begin{split}\dot{\mathcal{O}}=&-i[\mathcal{O},H_{\rm free}].\end{split}$ (3.8)
Additionally, we define $g_{\vec{k}}$ as
$\begin{split}g_{\vec{k}}(t)&\equiv\frac{1}{4V\omega_{k}}\sum_{r}\sum_{s}\left\langle b_{-\vec{k}}^{s}a_{\vec{k}}^{r}\right\rangle\bar{v}_{-\vec{k}}^{s}u_{\vec{k}}^{r},\\ g_{\vec{k}}^{*}(t)&\equiv\frac{1}{4V\omega_{k}}\sum_{r}\sum_{s}\left\langle a_{\vec{k}}^{r\dagger}b_{-\vec{k}}^{s\dagger}\right\rangle\bar{u}_{\vec{k}}^{r}v_{-\vec{k}}^{s},\end{split}$ (3.9)
then the time derivatives of $f_{\vec{k}}$ and $g_{\vec{k}}$ are given by
$\begin{split}\dot{f}_{\psi,\vec{k}}&=i\lambda\bar{\phi}\left(g_{\vec{k}}-g^{*}_{\vec{k}}\right)\cos m_{\phi}t,\\ \dot{f}_{\bar{\psi},\vec{k}}&=i\lambda\bar{\phi}\left(g_{-\vec{k}}-g^{*}_{-\vec{k}}\right)\cos m_{\phi}t,\\ \dot{g}_{\vec{k}}&=-i\left[2\omega_{k}g_{\vec{k}}(t)+\lambda\bar{\phi}\left\{\frac{k^{2}}{\omega_{k}^{2}}\left(1-f_{\psi,\vec{k}}-f_{\bar{\psi},-\vec{k}}\right)+\frac{2m_{\psi}}{\omega_{k}}g_{\vec{k}}\right\}\cos m_{\phi}t\right].\end{split}$ (3.10)
@@ 3.1 Without Cosmic Expansion
$\begin{split}\phi(t)=\bar{\phi}\cos m_{\phi}t.\end{split}$ (3.11)
@@
$\begin{split}g_{k}(t)=e^{-2i\omega_{k}t}[\xi_{k}(t)+i\eta_{k}(t)],\end{split}$ (3.12)
@@
$\begin{split}\tau=m_{\phi}t,\quad q=\lambda\bar{\phi}/m_{\phi},\quad\mathrm{and}\quad\tilde{\omega}_{k}=\omega_{k}/m_{\phi},\end{split}$ (3.13)
@@
$\begin{split}\frac{d{f}_{k}}{d\tau}&=q\left[2\xi_{k}\sin\left(2\tilde{\omega}_{k}\tau\right)-2\eta_{k}\cos\left(2\tilde{\omega}_{k}\tau\right)\right]\cos\tau,\\ \frac{d{\xi}_{k}}{d\tau}&=q\left[\left(1-2f_{k}\right)\sin(2\tilde{\omega}_{k}\tau)\right]\cos\tau,\\ \frac{d{\eta}_{k}}{d\tau}&=-q\left[\left(1-2f_{k}\right)\cos\left(2\tilde{\omega}_{k}\tau\right)\right]\cos\tau.\end{split}$ (3.14)
@@
$\begin{split}(1-2f_{k}(t))\dot{f_{k}}(t)=2\xi(t)\dot{\xi}(t)+2\eta(t)\dot{\eta}(t),\end{split}$ (3.15)
@@
$\begin{split}f_{k}(t)(1-f_{k}(t))=\xi_{k}^{2}(t)+\eta_{k}^{2}(t)=g_{k}(t)g_{k}^{*}(t)\geq 0,\end{split}$ (3.16)
@@
$\begin{split}\dot{\xi}_{k}(t)\cos(2\omega_{k}t)+\dot{\eta}_{k}(t)\sin(2\omega_{k}t)=0.\end{split}$ (3.17)

### v2

@@ 2.2 With Cosmic Expansion
In the expanding Universe, the magnitude of a physical momentum $k(t)\equiv|\vec{\bf k}(t)|$ redshifts inversely proportional to the scale factor in the metric of the spacetime. That is also true for the number density of the scalar field, in this case $n_{\phi}\propto a^{-3}$. Under these changes of the background, the scalar field $\phi$ can decay and fill the Fermi surface. However, the redshift of the momentum makes the Fermi surface empty and allows the decay of $\phi$. This continues until the number density of $\phi$ becomes smaller than the number density in the Fermi surface of momentum $p_{\psi}$, when Pauli blocking is not effective any more.
@@
$\begin{split}k(t_{p},t)=\frac{a(t_{p})}{a(t)}p_{\psi},\end{split}$ (22)
where $t_{p}$ is the production time of the fermion $\psi$ with momentum $p_{\psi}$. In matter-dominated Universe with $a(t)^{3}\propto t^{2}$, $k(t)$ has the relation to the time as
$\begin{split}t_{p}=t\left(\frac{|\vec{\bf k}(t)|}{p_{\psi}}\right)^{3/2}.\end{split}$ (23)
For a given comoving momentum, we can write the Boltzmann Eq. (10) in terms of a total derivative with respet to time as [11]
$\begin{split}\frac{df_{\psi}(t_{p},t)}{dt}=f_{\psi}^{\rm coll}(t_{p},t)\simeq\Gamma_{\phi}n_{\phi}(t)\left(\frac{\pi^{2}}{p_{\psi}^{2}}\right)\frac{\delta(t-t_{p})}{p_{\psi}H(t_{p})}\left[1-2f_{\psi}(\vec{\bf k}(t),t)\right],\end{split}$ (24)
@@
$\begin{split}\delta(k(t)-p_{\psi})=\frac{\delta(t-t_{p})}{p_{\psi}H(t_{p})}.\end{split}$ (25)
@@
$\begin{split}f_{\psi}(t_{p},t)=\frac{1}{2}\left(1-e^{-\bar{f}(t_{p})}\right)\theta(p_{\psi}-k(t)),\end{split}$ (26)
@@
$\begin{split}\bar{f}(t_{p})=\frac{\pi^{2}\Gamma_{\phi}}{p_{\psi}^{3}}\frac{n_{\phi}(t_{p})}{H(t_{p})}.\end{split}$ (27)
The Fermi surface of the momentum $k=p_{\psi}$, i.e. $f_{\psi}=1/2$, keeps being filled when the condition $\bar{f}(t_{p})\gg 1$ is still valid.
In Fig. 2 (left), we show the time evolution of the number density of scalar dark matter (solid line) and fermion field (dashed line) for different initial number density of scalar field with $10^{-30}\,\mathrm{GeV}$ (orange) and $n_{\phi}(t_{i})=10^{-25}\,\mathrm{GeV}$ (blue), respectively. For comparison, we also show the case of scalar decay without Pauli blocking (dotted line) for each case. When the number density of $\phi$ is large enough compared to that in the Fermi sphere,
$\begin{split}n_{\phi}(t)\gg\frac{p_{\psi}^{3}}{6\pi^{2}},\end{split}$ (28)
the fermions produced from the decay fills the fermi surface $k=p_{\psi}$ quickly with $f_{k}=1/2$ and the number density of them is
$\begin{split}n_{\psi}(t)=&g\int^{p_{\psi}}\frac{d^{3}k}{(2\pi)^{3}}f_{\psi}(k)=\frac{p_{\psi}^{3}}{6\pi^{2}},\end{split}$ (29)
@@
$\begin{split}\rho_{\psi}=\frac{p_{\psi}^{4}}{8\pi^{2}}.\end{split}$ (30)
@@
$\begin{split}\frac{d}{dt}(\rho_{\phi}+\rho_{\psi})+3H\rho_{\phi}+4H\rho_{\psi}=0.\end{split}$ (31)
@@
$\begin{split}\rho_{\phi}(t)=-\frac{p_{\psi}^{4}}{6\pi^{2}}+\left(\rho_{\phi}(t_{i})+\frac{p_{\psi}^{4}}{6\pi^{2}}\right)\left(\frac{a(t)}{a(t_{i})}\right)^{-3},\end{split}$ (32)
which is valid when $n_{\phi}(t)\gg p_{\psi}^{3}/6\pi^{2}$.
As the Universe expands, the number density of $\phi$ becomes small. When
$\begin{split}n_{\phi}(t)\lesssim\frac{p_{\psi}^{3}}{6\pi^{2}},\end{split}$ (33)
the number density of $\phi$ is not enough to fill the Fermi sphere and $\phi$ rapidly decays and $n_{\psi}$ begins to decrease only due to expansion of the Universe. Therefore, we can estimate the extended lifetime from the equality in Eq. (33), $n_{\phi}(t)=\frac{p_{\psi}^{3}}{6\pi^{2}}$. Using $n_{\phi}(t)\simeq n_{\phi}(t_{i})(a(t)/a_{i})^{-3}$, and $a(t)\propto t^{2/3}$ in matter-dominated era, we find that
$\begin{split}\tau_{\phi}\simeq t_{i}\left(\frac{\rho_{\phi}(t_{i})}{m_{\phi}}\right)^{1/2}\left(\frac{p_{\psi}^{3}}{4\pi^{2}}\right)^{-1/2}.\end{split}$ (34)
@@
$\begin{split}f_{\psi}(k,t_{0})=f_{\psi}\left(p_{\psi},t_{0}\left(\frac{k}{p_{\psi}}\right)^{3/2}\right).\end{split}$ (35)
@@
Figure 2: The time evolution of the number density of scalar (solid) and fermion (dashed) (left) and the distribution function of the produced fermions at momentum $p_{\psi}$ (right). Here, we used $m_{\phi}=10^{-3}\,{\rm eV},\ f_{\psi}(p_{\psi},t_{i})=0$ and $\Gamma_{\phi}=10^{-37}\,{\rm GeV}$, with $t_{0}=1/\Gamma_{\phi}$. The orange and blue lines correspond to the different initial number density, $n_{\phi}(t_{i})=10^{-30},10^{-25}\,\mathrm{GeV}^{3}$, respectively. The lifetime becomes longer for larger number density due to the Pauli blocking.
@@ 3 Particle Production with the Quantum Field Theory
$\begin{split}\psi(x)&=\sum_{s}\int\frac{d^{3}p}{(2\pi)^{3}}\frac{1}{\sqrt{2\omega_{p}}}\left(a_{\vec{p}}^{s}(t)u_{\vec{p}}^{s}e^{-i\vec{p}\cdot\vec{x}}+b_{\vec{p}}^{s\dagger}(t)v_{\vec{p}}^{s}e^{i\vec{p}\cdot\vec{x}}\right),\\ \bar{\psi}(x)&=\sum_{s}\int\frac{d^{3}p}{(2\pi)^{3}}\frac{1}{\sqrt{2\omega_{p}}}\left(a_{\vec{p}}^{s\dagger}(t)\bar{u}_{\vec{p}}^{s}e^{i\vec{p}\cdot\vec{x}}+b_{\vec{p}}^{s}(t)\bar{v}_{\vec{p}}^{s}e^{-i\vec{p}\cdot\vec{x}}\right),\end{split}$ (36)
where $\omega_{p}=\sqrt{m_{\psi}^{2}+|\vec{p}|^{2}}$ and the creation and annihilation operators satisfy the anti-commutation relations as
$\begin{split}\{a^{r}_{\vec{\bf k}},a^{s\dagger}_{\vec{\bf p}}\}=\{b^{r}_{\vec{\bf k}},b^{s\dagger}_{\vec{\bf p}}\}=(2\pi)^{3}\delta^{(3)}({\vec{\bf k}}-{\vec{\bf p}})\delta^{rs},\end{split}$ (37)
@@
$\begin{split}H_{\rm free}=\int\frac{d^{3}p}{(2\pi)^{3}}\sum_{s}\omega_{\vec{\bf p}}(a^{s\dagger}_{\vec{\bf p}}a^{s}_{\vec{\bf p}}+b^{s\dagger}_{\vec{\bf p}}b^{s}_{\vec{\bf p}}).\end{split}$ (38)
The expectation value of the operator $\mathcal{O}$ is defined by
$\begin{split}\expectationvalue{\mathcal{O}}\equiv{\rm Tr}(\rho\mathcal{O}),\end{split}$ (39)
using a density matrix $\rho$ which evolves as
$\rho(t)=U(t,t_{0})\rho(t_{0})U^{\dagger}(t,t_{0}),$ (40)
where $U(t,t_{0})$ is an unitary matrix. The evolution equation is given by
$i\frac{d}{dt}\rho(t)=\left[H,\rho\right],$ (41)
with $H$ being the total Hamiltonian, which can be divided into the free part and the interaction part as $H=H_{\mathrm{free}}+H_{\mathrm{int}}$. By substituting Eq. (41) into Eq. (39), we obtain
$\frac{d}{dt}\expectationvalue{\mathcal{O}}=$ $-i\expectationvalue{\commutator{\mathcal{O}}{H_{\rm int}}}+\dot{\expectationvalue{\mathcal{O}}},$ , (42)
$\dot{\mathcal{O}}=$ $-i[\mathcal{O},H_{\rm free}].$ (43)
Applying Eq. (42) with the interaction Hamiltonian (1) to the operators of fermions $a_{\vec{p}}^{\dagger}a_{\vec{p}}$, $b_{\vec{p}}^{\dagger}b_{\vec{p}}$, and $b_{\vec{p}}a_{\vec{p}}$, we obtain differential equations
$\begin{split}\dot{f}_{\psi,\vec{k}}&=i\lambda\phi_{c}(t)\left(g_{\vec{k}}-g^{*}_{\vec{k}}\right),\\ \dot{f}_{\bar{\psi},\vec{k}}&=i\lambda\phi_{c}(t)\left(g_{-\vec{k}}-g^{*}_{-\vec{k}}\right),\\ \dot{g}_{\vec{k}}&=-i\left[2\omega_{k}g_{\vec{k}}(t)+\lambda\phi_{c}(t)\left\{\frac{k^{2}}{\omega_{k}^{2}}\left(1-f_{\psi,\vec{k}}-f_{\bar{\psi},-\vec{k}}\right)+\frac{2m_{\psi}}{\omega_{k}}g_{\vec{k}}\right\}\right],\end{split}$ (44)
where $f_{\vec{k}}$ and $g_{\vec{k}}$ are defined by
$f_{\psi,\vec{k}}(t)$ $\equiv\frac{1}{2V}\sum_{r}\left\langle a_{\vec{k}}^{r\dagger}a_{\vec{k}}^{r}\right\rangle,$ (45a)
$f_{\bar{\psi},\vec{k}}(t)$ $\equiv\frac{1}{2V}\sum_{r}\left\langle b_{\vec{k}}^{r\dagger}b_{\vec{k}}^{r}\right\rangle,$ (45b)
$g_{\vec{k}}(t)$ $\equiv\frac{1}{4V\omega_{k}}\sum_{r}\sum_{s}\left\langle b_{-\vec{k}}^{s}a_{\vec{k}}^{r}\right\rangle\bar{v}_{-\vec{k}}^{s}u_{\vec{k}}^{r},$ (45c)
$g_{\vec{k}}^{*}(t)$ $\equiv\frac{1}{4V\omega_{k}}\sum_{r}\sum_{s}\left\langle a_{\vec{k}}^{r\dagger}b_{-\vec{k}}^{s\dagger}\right\rangle\bar{u}_{\vec{k}}^{r}v_{-\vec{k}}^{s}.$ (45d)
We note that the number density operator $n_{\vec{p}}$ in the density matrix formalism is related in terms of the distribution function $f_{\vec{p}}$ as
$n=\int\frac{d^{3}p}{(2\pi)^{3}}f_{\vec{p}}.$ (46)
Since we consider that fermions are produced only from the decay of scalar field, initial quantum states are taken to be the vacuum state $|0\rangle$ so that the density operator is $\rho(0)=|0\rangle\langle 0|$ which gives the initial condition
$\begin{split}f_{\psi,\vec{k}}(0)=f_{\bar{\psi},\vec{k}}(0)=g_{\vec{k}}(0)=0.\end{split}$ (47)
@@
In principle, we could also consider the distribution function of the scalar field in terms of the creation and annihilation operators and derive the evolution equation, which enables us to follow the number density of the scalar field when it starts to decay. However, in this study we focus on the stability of the scalar particle and only show that the production of the fermions are restricted due to the Pauli blocking. The full treatment might be beyond the present study and will be done in another publication.
In the following, we use an approximation that the scalar field is dominated by the classical field $\phi(x)\simeq\phi_{c}(t)$. The quantum particles of the scalar field might be produced from the back-reaction of $\psi\bar{\psi}\to\phi$. However, since we will focus only on the timescales where the number density of the fermions are negligible compared to that of the classical scalar field, the number density of the quantum scalar particle also can be ignored in our consideration.
@@ 3.1 Without Cosmic Expansion
$\begin{split}\phi_{c}(t)=\bar{\phi}\cos m_{\phi}t.\end{split}$ (48)
@@
$\begin{split}g_{k}(t)=e^{-2i\omega_{k}t}[\xi_{k}(t)+i\eta_{k}(t)],\end{split}$ (49)
@@
$\begin{split}\tau=m_{\phi}t,\quad q=\lambda\bar{\phi}/m_{\phi},\quad\mathrm{and}\quad\tilde{\omega}_{k}=\omega_{k}/m_{\phi},\end{split}$ (50)
@@
$\begin{split}\frac{d{f}_{k}}{d\tau}&=q\left[2\xi_{k}\sin\left(2\tilde{\omega}_{k}\tau\right)-2\eta_{k}\cos\left(2\tilde{\omega}_{k}\tau\right)\right]\cos\tau,\\ \frac{d{\xi}_{k}}{d\tau}&=q\left[\left(1-2f_{k}\right)\sin(2\tilde{\omega}_{k}\tau)\right]\cos\tau,\\ \frac{d{\eta}_{k}}{d\tau}&=-q\left[\left(1-2f_{k}\right)\cos\left(2\tilde{\omega}_{k}\tau\right)\right]\cos\tau.\end{split}$ (51)
@@
$\begin{split}(1-2f_{k}(t))\dot{f_{k}}(t)=2\xi(t)\dot{\xi}(t)+2\eta(t)\dot{\eta}(t),\end{split}$ (52)
@@
$\begin{split}f_{k}(t)(1-f_{k}(t))=\xi_{k}^{2}(t)+\eta_{k}^{2}(t)=g_{k}(t)g_{k}^{*}(t)\geq 0,\end{split}$ (53)
@@
$\begin{split}\dot{\xi}_{k}(t)\cos(2\omega_{k}t)+\dot{\eta}_{k}(t)\sin(2\omega_{k}t)=0.\end{split}$ (54)